100 - T = 125.. My daughter who is in 4th grade brought this problem home on her math homework page and I cannot remember this far back!! Sounds horrible, I know. HELP PLEASE! :0)
Answer by ankor@dixie-net.com(22740)

100 - T = 125
Subtract 100 from both sides:
-T = 125 - 100
-T = 25
The variable (T in this case) has to be positive, multiply both side by -1
T = -25
:
:
Check solution in original equation
100 - (-25) = 120
Remember that minus a minus is a plus, so you have:
100 + 25 = 125; confirms our solution